OpenStudy (anonymous):

Mrs. Perrin, the school music teacher, spent twice as much money on sheet music as she did on CDs. The total amount of money that she spent for sheet music and CDs was one third of her monthly music budget of $189. How much money did Mrs. Perrin spend on sheet music?

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) from this problem you can get two equations (sheet music cost) = 2*(cd cost) (sheet music + cd) = (1/3)*(budget)